Flip a fair coin 4 times. What is the probability of observong exactly three heads?
*Should I use binomial distribution or HHH method?

Answer by stanbon(75887)

Flip a fair coin 4 times. What is the probability of observong exactly three heads?
*Should I use binomial distribution or HHH method?
Yes
-----------------
Binomial Problems with n = 4 ; p(H) = 1/2
------------------------
P(x = 3) = 4C3(1/2)^3*(1/2) = 4*(1/2)4 = 4/16 = 1/4
--------------------------------------------------------
Cheers,
Stan H.
=============